# QMol-grid: A MATLAB package for quantum-mechanical simulations in atomic and molecular systems

### v2

@@ Abstract
The QMol-grid package provides a suite of routines for performing quantum-mechanical simulations in atomic and molecular systems in one spatial dimension. It supports ground- and excited-state calculations for the Schrödinger equation, density-functional theory, and Hartree-Fock levels of theory as well as propagators for field-free and field-driven time-dependent Schrödinger equation (TDSE) and real-time time-dependent density-functional theory (TDDFT), using symplectic-split schemes. The package is written using MATLAB’s object-oriented features and handle classes. It is designed to facilitate access to the wave function(s) (TDSE) and the Kohn-Sham orbitals (TDDFT) within MATLAB’s environment.
@@
Table 1: QMol-grid metadata
@@ 1 Motivation and significance
Ab initio quantum simulations of the electronic structure and dynamics in atoms and molecules play an important role in many fields of physics and chemistry. They have lead to the development of many computational packages. For instance, optimized packages like [2, 3, 4, 5, 6] allow for routine quantum calculations in a range of atomic, molecular, and solid-state systems, typically running on high-performance computer (HPC) systems. Alternatively, the QMol-grid package has been developed in the context of ultrafast atomic, molecular, and optical (AMO) research [7, 8] (i) to provide a test bed for quantum-mechanical simulations that can easily run on personal computers, including when considering molecular systems with multiple interacting electrons and (ii) to facilitate access to the wave function(s) (TDSE) and the Kohn-Sham orbitals (TDDFT), such that users can build complex workflows and analyses alongside the simulations. For instance, QMol-grid time propagators enable arbitrary user-defined functions to be evaluated, and their result stored, while the TDSE/TDDFT propagation is performed. The package also provides built-in facilities for the calculation of common observable, including the dipole signal, energy, ionization, TDSE wave function and TDDFT Kohn-Sham orbitals. Aside from research purposes, the package offers a valuable resource for teaching purposes: with it, students can be introduced to a range of quantum mechanical simulation techniques (see below), using calculation examples that run on personal computers or laptops.
The QMol-grid package provides a suite of routines for performing quantum-mechanical simulations in atomic and molecular systems in one spatial dimension. All simulations use an underlying Cartesian-grid discretization scheme, with all spatial derivatives calculated with fast-Fourier transforms. The package is written using MATLAB’s object-oriented features and handle classes. Notably, the package supports:
@@
We refer readers to the documentation for details regarding each supported computational framework. Briefly, within QMol-grid, SE provides a single-active electron model of the electronic structure of atoms and molecules. For multi-electron systems, HF gives the best approximation (lowest energy) of the wave function in terms of a single antisymmetrized product of one-electron wave functions (Slater determinant) [9]. Alternatively, DFT trades the multi-electron wave-function picture for the real-space electron density, whose dimension is independent of the number of active electrons. There, electron-electron interactions are captured in the (nonlinear) functional dependency of the DFT Hamiltonian on the electron density. Specifically, QMol-grid uses Kohn-Sham DFT [10], where the density is build from virtually-independent electrons. Both HF and DFT correspond to solving a nonlinear eigen-state problem, which is implemented via standard iterative techniques in the package [11].
TDSE and TDDFT describe the time evolution of the system, typically either resulting from an external driving laser field or starting from a non-stationary initial state, within their respective SE and DFT framework. From its origin in ultrafast AMO science research, the QMol-grid package offers efficient and high-order time propagation schemes specially designed for those simulations [8]. Time-dependent simulations neglect nuclear dynamics (Born-Oppenheimer approximation), with all atomic and molecular potentials fixed in space throughout the time evolution of the electrons.
@@ 2 Software description
A full description of the QMol-grid package, including all possible input parameters and calculation features is included in the MATLAB documentation provided with the package. After installation, the package documentation is accessible in MATLAB, in the “Supplemental Software” section. A copy of the documentation is also provided on the GitHub wiki. The documentation includes a series of tutorials, starting with SE ground-state calculations, and going through TDSE, DFT, and TDDFT calculations to help new users getting familiarized with setting up calculations, input parameters, and output variables. Throughout, the documentation also includes many script samples illustrating how one can use the various features. Finally, the documentation discusses the required class structure for advanced users who wish to add their own functionalities to the package and inherit common interface methods to the QMol-grid package.
@@ 2.1 Software architecture
The QMol-grid package provides an ecosystem of MATLAB handle classes. While the package is provided as a stand-alone suite, it is developed around 3 main groups sketched in figure 1 (a): (1) external components, (2) kernel classes that define high-level calculation methods, and (3) implementation classes that define all the lower-level functionalities. The package is developed with the general goal of facilitating access to the wave function(s) (SE/TDSE) and the Kohn-Sham orbitals (DFT/TDDFT), which are packaged into classes for abstract manipulations of the objects in ground-state, time propagation, and common observables’ calculations.
Figure 1: (a) Overall architecture for the QMol-grid package: components are sorted in three tiers of handle classes that define the computation ecosystem. (b) Schematic of a Schrödinger-equation object and its components. Each box indicates a separate class defined within the package.
Users set up calculations by creating QMol-grid objects of the relevant type and specifying the desired parameters using MATLAB’s common name-value pair argument structure (in arbitrary order and case insensitive). As illustrated in the examples of section 3 below, we strive to give intuitive and descriptive parameter names. The documentation provides the list, together with supported formats, of all available input parameters for each class. Throughout the package, input parameters and output results are specified in atomic units; we provide units conversion external components to facilitate conversions to more conventional units (e.g., as/fs for time, W/cm2 for field intensity, etc.). Some high-level components are themselves encapsulated into classes, enabling abstract manipulations in the property objects. Figure 1 (b) illustrates this concept for the Schrödinger-equation object QMol_SE for which class properties are a mix of variables (xspan and numberWaveFunction) and QMol-grid objects (discretization, waveFunction, and potential). Parameters can be updated after an object has been created using the set method, again using name-value pair arguments.
@@ 2.2 Software functionalities
Ground- and excited-state calculations in the QMol-grid package are performed by a direct diagonalization of the Hamiltonian operator, via MATLAB’s eigs (grid discretization) or eig (basis set) functions. DFT and HF self-consistent-field iterations are performed using an Anderson’s mixing scheme [12, 11]. HF is obtained by running DFT with an exact-exchange and no correlation functionals.
@@
    Checkpointing, with the creation of a restart MATLAB file (.mat) that can be used to resume a calculation that was stopped before it was finished;
@@
    Saving the intermediate Schrödinger- or DFT-model objects in separate MATLAB files (.mat).
Aside from the options that generate MATLAB files (first and last items above), the results for all the other on-the-fly calculations are collected and stored in the time propagator object itself – see the TDDFT example in section 3.2. The size of the generated output strongly depends on the simulation parameters: time-dependent dipole, energy, and ionization signals are proportional to the number of saved time steps while wave functions, Kohn-Sham orbitals, and densities scale as the number of time steps multiplied by the domain grid size. Anecdotally, in our experience dipole, energy, and ionization signals typically require a few hundred KB while saving the wave function or density easily takes a few to many MB.
@@
The QMol-grid package comes with a suite of unit tests, individually checking the methods in each of the classes in the package.
@@ 3 Illustrative examples
We illustrate how users interface with the QMol-grid package in two examples. The documentation includes a more comprehensive series of tutorials meant to get new users familiarized with how to set simulations up, interact, and recover results from calculations. Starting from ground-state SE and moving towards TDDFT, the tutorials progressively introduce (i) minimal-code examples and (ii) discussions of various input parameters and output variables available in the package.
@@ 3.1 Example 1: Schrödinger-equation ground state
Here we illustrate how to use the QMol-grid package to calculate the ground-state wave function of a one-dimensional hydrogen-like atom. The Schrödinger-equation ground-state corresponds to the lowest-energy solution to the eigenvalue problem ^$\hat{\mathcal{H}}\psi(x)=E\psi(x)$, where $\hat{\mathcal{H}}$ is the Schrödinger-equation Hamiltonian operator, $\psi$ is the wave function, and $E$ its associated energy. In atomic units, the Hamiltonian operator is ^$\hat{\mathcal{H}}=-\frac{\Delta}{2}+\hat{\mathcal{V}}$.
@@
where ’softeningParameter’ specifies the value for the parameter $a$. Here we choose the softening parameter $a=\sqrt{2}$ to match H’s ground state energy. By default, the atom is located at the origin $x=0$. Note that H only corresponds to the atomic model, which is shared with molecular systems and various quantum frameworks. Thus, it must be turned into a valid Schrödinger-equation potential, using
@@
Here ’atom’ indicates to the QMol_SE_V object that the list of atomic centers is provided next – here a single H effective potential.
@@
Like above, when creating the SE object, we recognize the definition of the discretization domain and effective potential with the keywords ’xspan’ and ’potential’, respectively. Next we move to calculating its associated ground-state wave function and energy using the two commands
@@
The first line creates the eigen-state solver while the second performs the actual ground-state calculation on the Schrödinger-equation object SE. At the end of the calculation, the ground-state wave function is stored in the input SE, together with relevant information such as the domain discretization. For instance, solely relying on SE, one can plot the ground-state wave function with
@@
The output is represented in Fig. 2. From the plot command line, we see that the domain-discretization grid may be recovered using the xspan property in the object SE (using the standard object-oriented dot notation SE.xspan). On the other hand, the wave function is nested inside another object, which explains the consecutive dots SE.waveFunction.waveFunction. Other properties in the object SE.waveFunction are used by ground/excited-state and TDSE calculations; we refer to the QMol_SE_wfcn documentation page for further details.
@@ 3.2 Example 2: Time-dependent density-functional theory
The QMol-grid package relies on the canonical Hamiltonian structure of TDDFT [8] to integrate the dynamics of equation (1). In this example, we illustrate how to use the QMol-grid package to integrate the TDDFT dynamics of an open-shell one-dimensional molecular ion model with 3 atomic centers and 5 active electrons.
Initial condition: In the QMol-grid package, TDDFT simulations are decoupled from setting up the initial condition, which must be done independently. Similar to example 1, we build the molecular model out of 3 one-dimensional atomic models, each contributing 2 electrons to the molecule, using soft-Coulomb potentials. For our example, we start by calculating the neutral-molecule ground state:
@@
    V_1     =   QMol_Va_softCoulomb(’atom’,’X_1’,’charge’,2,’position’,-3);
    V_2     =   QMol_Va_softCoulomb(’atom’,’X_2’,’charge’,2,’position’, 0);
    V_3     =   QMol_Va_softCoulomb(’atom’,’X_3’,’charge’,2,’position’, 3);
@@
The “Molecular model” block defines the atomic effective potential, specifying the name, bare charge, and location of each atomic center, respectively. The “DFT model” block first defines the molecular potential Vext, followed by the DFT functionals Vh and Vxc to be used in the (TD)DFT calculations – see the documentation’s ground-state DFT tutorial for further details regarding the model parameters. The final block “DFT ground state” first creates the eigen-state DFT solver, here an Anderson mixing scheme [11], and performs the ground-state self-consistent field (SCF) calculation.
@@
Plotting the result: Next we recover calculated observables out of the TDDFT object. Each set of observable is stored in a separate structure property in the TDDFT object, which contains (i) the exact time vector at which the quantity has been saved and (ii) the observable itself. In our case, the structure of interest is TDDFT.outDensity with the up- and down-spin densities respectively stored in the fields totalUp and totalDown. The densities are matrices with columns corresponding to the successive saved times. To plot the spin density, defined as the difference between the up- and down-spin one-body densities, we use
@@
Profiling (estimating the memory footprint): Before running the TDDFT calculation, users have the possibility to check how much memory the simulation requires to run and store the requested one-body densities. Using the same calculation workflow as above, right after creating the TDDFT object, the memory footprint is obtained with
@@
In our case, the estimated total TDDFT-object size is 1.8 MB with 1.5 MB for the saved electron density. Saving the TDDFT and DFT object in a MATLAB file at the end of the propagation produces a 1.6 MB .mat file. We mostly attribute the slight difference with the profiler estimate to run-time memory overhead associated with internal variables that are not stored in the saved objects.

### v3

# |QMol-grid|:A MATLAB package for quantum-mechanical simulations in atomic and molecular systems

@@ Abstract
The |QMol-grid|package provides a suite of routines for performing quantum-mechanical simulations in atomic and molecular systems, currently implemented in one spatial dimension. It supports ground- and excited-state calculations for the Schrödinger equation, density-functional theory, and Hartree-Fock levels of theory as well as propagators for field-free and field-driven time-dependent Schrödinger equation (TDSE) and real-time time-dependent density-functional theory (TDDFT), using symplectic-split schemes. The package is written using MATLAB’s object-oriented features and handle classes. It is designed to facilitate access to the wave function(s) (TDSE) and the Kohn-Sham orbitals (TDDFT) within MATLAB’s environment.
@@
Table 1: |QMol-grid|metadata
@@ 1 Motivation and significance
Ab initio quantum simulations of the electronic structure and dynamics in atoms and molecules play an important role in many fields of physics and chemistry. They have lead to the development of many computational packages. For instance, optimized packages like [2, 3, 4, 5, 6] allow for routine quantum calculations in a range of atomic, molecular, and solid-state systems, typically running on high-performance computer (HPC) systems. Alternatively, the |QMol-grid|package has been developed in the context of ultrafast atomic, molecular, and optical (AMO) research [7, 8], with a focus on low-dimension atomic and molecular models, (i) to provide a test bed for quantum-mechanical simulations that can easily run on personal computers, including when considering molecular systems with multiple interacting electrons and (ii) to facilitate access to the wave function(s) (TDSE) and the Kohn-Sham orbitals (TDDFT), such that users can build complex workflows and analyses alongside the simulations. For instance, |QMol-grid|time propagators enable arbitrary user-defined functions to be evaluated, and their result stored, while the TDSE/TDDFT propagation is performed. The package also provides built-in facilities for the calculation of common observable, including the dipole signal, energy, ionization, TDSE wave function and TDDFT Kohn-Sham orbitals. Aside from research purposes, the package offers a valuable resource for teaching purposes: with it, students can be introduced to a range of quantum mechanical simulation techniques (see below), using calculation examples that run on personal computers or laptops.
The |QMol-grid|package provides a suite of routines for performing quantum-mechanical simulations in atomic and molecular systems, currently implemented in one spatial dimension. Obviously, such lower-dimensional models cannot capture the entire manifold of processes at play in full-dimension simulations. Instead, these models play an important and complementary role in providing prototypical systems where general, non-system specific, properties can be established. A second advantage of dimensionally-reduced simulations is that they typically run at a fraction of the time of their full-dimension counterparts. This computational up-speed can then be re-invested in extended parameter scans or scouting for outcome of interest in a large parameter space. For instance, we have used this latter approach in recent analyses of ultrafast migration of charges in molecules [7]. The specifics of what is included and left out in any given lower-dimension simulation is highly system/model dependent. We defer to end-users of the package to address those limitations in their specific situation.
All simulations in the |QMol-grid|package use an underlying Cartesian-grid discretization scheme, with all spatial derivatives calculated with fast-Fourier transforms. The package is written using MATLAB’s object-oriented features and handle classes. Notably, the package supports:
@@
We refer readers to the documentation for details regarding each supported computational framework. Briefly, within |QMol-grid|,SE provides a single-active electron model of the electronic structure of atoms and molecules. For multi-electron systems, HF gives the best approximation (lowest energy) of the wave function in terms of a single antisymmetrized product of one-electron wave functions (Slater determinant) [9]. Alternatively, DFT trades the multi-electron wave-function picture for the real-space electron density, whose dimension is independent of the number of active electrons. There, electron-electron interactions are captured in the (nonlinear) functional dependency of the DFT Hamiltonian on the electron density. Specifically, |QMol-grid|uses Kohn-Sham DFT [10], where the density is build from virtually-independent electrons. Both HF and DFT correspond to solving a nonlinear eigen-state problem, which is implemented via standard iterative techniques in the package [11].
TDSE and TDDFT describe the time evolution of the system, typically either resulting from an external driving laser field or starting from a non-stationary initial state, within their respective SE and DFT framework. From its origin in ultrafast AMO science research, the |QMol-grid|package offers efficient and high-order time propagation schemes specially designed for those simulations [8]. Time-dependent simulations neglect nuclear dynamics (Born-Oppenheimer approximation), with all atomic and molecular potentials fixed in space throughout the time evolution of the electrons.
@@ 2 Software description
A full description of the |QMol-grid|package, including all possible input parameters and calculation features is included in the MATLAB documentation provided with the package. After installation, the package documentation is accessible in MATLAB, in the “Supplemental Software” section. A copy of the documentation is also provided on the GitHub wiki. The documentation includes a series of tutorials, starting with SE ground-state calculations, and going through TDSE, DFT, and TDDFT calculations to help new users getting familiarized with setting up calculations, input parameters, and output variables. Throughout, the documentation also includes many script samples illustrating how one can use the various features. Finally, the documentation discusses the required class structure for advanced users who wish to add their own functionalities to the package and inherit common interface methods to the |QMol-grid|package.
@@ 2.1 Software architecture
The |QMol-grid|package provides an ecosystem of MATLAB handle classes. While the package is provided as a stand-alone suite, it is developed around 3 main groups sketched in figure 1 (a): (1) external components, (2) kernel classes that define high-level calculation methods, and (3) implementation classes that define all the lower-level functionalities. The package is developed with the general goal of facilitating access to the wave function(s) (SE/TDSE) and the Kohn-Sham orbitals (DFT/TDDFT), which are packaged into classes for abstract manipulations of the objects in ground-state, time propagation, and common observables’ calculations.
Figure 1: (a) Overall architecture for the |QMol-grid|package: components are sorted in three tiers of handle classes that define the computation ecosystem. (b) Schematic of a Schrödinger-equation object and its components. Each box indicates a separate class defined within the package.
Users set up calculations by creating |QMol-grid|objects of the relevant type and specifying the desired parameters using MATLAB’s common name-value pair argument structure (in arbitrary order and case insensitive). As illustrated in the examples of section 3 below, we strive to give intuitive and descriptive parameter names. The documentation provides the list, together with supported formats, of all available input parameters for each class. Throughout the package, input parameters and output results are specified in atomic units; we provide units conversion external components to facilitate conversions to more conventional units (e.g., as/fs for time, W/cm2 for field intensity, etc.). Some high-level components are themselves encapsulated into classes, enabling abstract manipulations in the property objects. Figure 1 (b) illustrates this concept for the Schrödinger-equation object |QMol_SE|for which class properties are a mix of variables (|xspan|and |numberWaveFunction|)and |QMol-grid|objects (|discretization|,|waveFunction|,and |potential|). % Parameterscan be updated after an object has been created using the |set|method, again using name-value pair arguments.
@@ 2.2 Software functionalities
Ground- and excited-state calculations in the |QMol-grid|package are performed by a direct diagonalization of the Hamiltonian operator, via MATLAB’s |eigs|(grid discretization) or |eig|(basis set) functions. DFT and HF self-consistent-field iterations are performed using an Anderson’s mixing scheme [12, 11]. HF is obtained by running DFT with an exact-exchange and no correlation functionals.
@@
    Checkpointing, with the creation of a restart MATLAB file (|.mat|)that can be used to resume a calculation that was stopped before it was finished;
@@
[…]
@@
    The |QMol-grid|package comes with a suite of unit tests, individually checking the methods in each of the classes in the package.
@@ 3 Illustrative examples
    We illustrate how users interface with the |QMol-grid|package in two examples. The documentation includes a more comprehensive series of tutorials meant to get new users familiarized with how to set simulations up, interact, and recover results from calculations. Starting from ground-state SE and moving towards TDDFT, the tutorials progressively introduce (i) minimal-code examples and (ii) discussions of various input parameters and output variables available in the package.
@@ 3.1 Example 1: Schrödinger-equation ground state
    Here we illustrate how to use the |QMol-grid|package to calculate the ground-state wave function of a one-dimensional hydrogen-like atom. The Schrödinger-equation ground-state corresponds to the lowest-energy solution to the eigenvalue problem ^$\hat{\mathcal{H}}\psi(x)=E\psi(x)$, where $\hat{\mathcal{H}}$ is the Schrödinger-equation Hamiltonian operator, $\psi$ is the wave function, and $E$ its associated energy. In atomic units, the Hamiltonian operator is ^$\hat{\mathcal{H}}=-\frac{\Delta}{2}+\hat{\mathcal{V}}$.
@@
    where |’softeningParameter’|specifies the value for the parameter $a$. Here we choose the softening parameter $a=\sqrt{2}$ to match H’s ground state energy. By default, the atom is located at the origin $x=0$. Note that H only corresponds to the atomic model, which is shared with molecular systems and various quantum frameworks. Thus, it must be turned into a valid Schrödinger-equation potential, using
@@
    Here |’atom’|indicates to the |QMol_SE_V|object that the list of atomic centers is provided next – here a single H effective potential.
@@
    Like above, when creating the |SE|object, we recognize the definition of the discretization domain and effective potential with the keywords |’xspan’|and |’potential’|,respectively. Next we move to calculating its associated ground-state wave function and energy using the two commands
@@
    The first line creates the eigen-state solver while the second performs the actual ground-state calculation on the Schrödinger-equation object |SE|. % Atthe end of the calculation, the ground-state wave function is stored in the input |SE|,together with relevant information such as the domain discretization. For instance, solely relying on |SE|,one can plot the ground-state wave function with
@@
    The output is represented in Fig. 2. From the plot command line, we see that the domain-discretization grid may be recovered using the |xspan|property in the object |SE|(using the standard object-oriented dot notation |SE.xspan|).On the other hand, the wave function is nested inside another object, which explains the consecutive dots |SE.waveFunction.waveFunction|. % Otherproperties in the object |SE.waveFunction|are used by ground/excited-state and TDSE calculations; we refer to the |QMol_SE_wfcn|documentation page for further details.
@@ 3.2 Example 2: Time-dependent density-functional theory
    The |QMol-grid|package relies on the canonical Hamiltonian structure of TDDFT [8] to integrate the dynamics of equation (1). In this example, we illustrate how to use the |QMol-grid|package to integrate the TDDFT dynamics of an open-shell one-dimensional molecular ion model with 3 atomic centers and 5 active electrons.
    Initial condition: In the |QMol-grid|package, TDDFT simulations are decoupled from setting up the initial condition, which must be done independently. Similar to example 1, we build the molecular model out of 3 one-dimensional atomic models, each contributing 2 electrons to the molecule, using soft-Coulomb potentials. For our example, we start by calculating the neutral-molecule ground state:
@@
    V_1 = QMol_Va_softCoulomb( ...
    ’atom’,’X_1’,’charge’,2,’position’,-3);
    V_2 = QMol_Va_softCoulomb( ...
    ’atom’,’X_2’,’charge’,2,’position’, 0);
    V_3 = QMol_Va_softCoulomb( ...
    ’atom’,’X_3’,’charge’,2,’position’, 3);
@@
    The “|%Molecular model—” block defines the atomic effective potential, specifying the name, bare charge, and location of each atomic center, respectively. The “|%DFT model—” block first defines the molecular potential |Vext|,followed by the DFT functionals |Vh|and |Vxc|to be used in the (TD)DFT calculations – see the documentation’s ground-state DFT tutorial for further details regarding the model parameters. The final block “|%DFT ground state—” first creates the eigen-state DFT solver, here an Anderson mixing scheme [11], and performs the ground-state self-consistent field (SCF) calculation.
@@
    Plotting the result: Next we recover calculated observables out of the TDDFT object. Each set of observable is stored in a separate structure property in the TDDFT object, which contains (i) the exact time vector at which the quantity has been saved and (ii) the observable itself. In our case, the structure of interest is |TDDFT.outDensity|with the up- and down-spin densities respectively stored in the fields |totalUp|and |totalDown|.The densities are matrices with columns corresponding to the successive saved times. To plot the spin density, defined as the difference between the up- and down-spin one-body densities, we use
@@
    Profiling (estimating the memory footprint): Before running the TDDFT calculation, users have the possibility to check how much memory the simulation requires to run and store the requested one-body densities. Using the same calculation workflow as above, right after creating the |TDDFT|object, the memory footprint is obtained with
@@
    In our case, the estimated total TDDFT-object size is 1.8 MB with 1.5 MB for the saved electron density. Saving the |TDDFT|and |DFT|object in a MATLAB file at the end of the propagation produces a 1.6 MB |.mat|file. We mostly attribute the slight difference with the profiler estimate to run-time memory overhead associated with internal variables that are not stored in the saved objects.